OpenStudy (anonymous):

log base 3 (x-9) + log base 3 (x-3) = 2

OpenStudy (anonymous):

a law of logs says that \[\log_{a} b +\log_{a} c =\log_{a}bc \] try now

OpenStudy (anonymous):

log base 3 (x-9)(x-3)=2

OpenStudy (anonymous):

base3 xsquared -12x + 27 = 2

OpenStudy (raden):

hint : setting in Right side be log base3 9

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

yeah then you can change that to \[3^{2}=x ^{2}-12x+27\]

OpenStudy (anonymous):

another rule of logs

OpenStudy (raden):

2 = log base3 9, right ? because 3^2 = 9

OpenStudy (anonymous):

x^2-12x=-18

OpenStudy (anonymous):

hint: form a quadratic to use the quadratic formula